OpenStudy (anonymous):

Solve by subsitution y= 2/3x-2 2x+3y=9

OpenStudy (anonymous):

\[2x+3(\frac{2}{3}x -2) = 9\] \[2x+2x -6 = 9\] \[4x -6 = 9\] \[4x = 15\] \[x = \frac{15}{4}\] plug it back into the first eq. \[y = \frac{2}{3} \cdot \frac{15}{4} -2 \] \[y = \frac{30}{12} -2 \] \[y = \frac{5}{2} -\frac{4}{2} \] \[y = \frac{1}{2} \]